# Large Language Model Can Continue Evolving From Mistakes

### v5

@@ 5.1 Main Results
Firstly, Extractive and Review Corpus significantly enhance LLM performance. CEM-E consistently outperforms its control group, showing higher accuracy and W2R metrics. Similarly, CEM-R outperforms its control group and significantly reduces the R2W metric.
The results indicate an improvement in the W2R metric due to the use of the Extractive Corpus, suggesting an enhanced capacity of the model to identify and comprehend relevant knowledge within the CPT data. This improvement leads to more effective error correction. Additionally, the Review Corpus significantly reduces the R2W metric, indicating a suppression of the model’s tendency to forget correct knowledge.
Secondly, a Supplemental Corpus from multiple sources does not invariably enhance the CEM method more than a single source. Results in Figure 4 indicate no substantial difference in CEM effectiveness, whether using Wikipedia, Bing, or their combination. Notably, RMix records lower accuracy compared to RWiki and RBing, despite a larger dataset. This may arise from data redundancies. Moreover, an excessive supplemental corpus can reduce the proportion of CIT data, undermining its role in mitigating catastrophic forgetting, improving knowledge extraction, and increasing the risk of overfitting.
Thirdly, proportion factor $\boldsymbol{\alpha}$ has limited impact on the effectiveness of Review Corpus. Table 9 shows that R0, R0.5, and R1 have similar effects on LLM performance. This suggests that reviewing the knowledge in any format consolidates it. Therefore, unless specified, R will refer to the R0.5 in the following.
@@ 5.2 Multiple Iterations Analysis
The CEM method consistently yields significant improvements across multiple rounds, underscoring its iterative potential. We conduct a second round of CEM on CuteGPT-13B for the Xiezhi task, employing the same training strategies and data sources as in the first round. Independent CIT data of equivalent volume are used. Results in Table 10 indicate that this round significantly enhances the LLM’s accuracy across all dataset construction strategies. Notably, RWiki shows a 4.75% increase in accuracy over the previous round and a 17.00% improvement over INIT.
Further analysis shows that the Review Strategy consistently has significantly lower R2W metrics compared to other dataset construction strategies. For example, the R2W metric of RWiki decreases by 7.02% compared to PWiki and by 5.90% compared to EWiki. We argue that the Review Strategy significantly reduces the probability of changing a previously correct response to an incorrect one, thereby helping the model avoid severe forgetting during multiple rounds of supplemental training.
Figure 5: The figure shows the experimental results of CuteGPT-13B on the Xiezhi task after multiple rounds of CEM. The left part represents results without replay, while the right part shows results with a random replay strategy. ER indicates the enhancement rate on the Xiezhi task, and AFR represents the average forgetting rate on the HotpotQA and GSM8K tasks.
Catastrophic Forgetting Analysis
@@
To analyze catastrophic forgetting, we test CuteGPT-13B on Xiezhi, HotpotQA, and GSM8K to evaluate domain knowledge, reasoning, and mathematical skills. We use the CEM method to improve performance on Xiezhi and monitor any declines in the other tasks. Initially, we sample 2,000 instances from each dataset to establish a baseline. To mitigate forgetting, we employ random replay by replacing the General Corpus with 500 randomly replayed instances from HotpotQA and GSM8K in each training round.
@@
Figure 5 shows that even after three rounds of CEM, performance on Xiezhi improves significantly and maintains an upward trend. The combination of CEM and random replay effectively mitigates performance decline, with AFR exceeding 70.0% in most cases. However, without a replay strategy, initial rounds show better improvement (possibly because the model does not need to adapt to the more complex multitasking required by the random replay strategy), but more rounds lead to greater forgetting, limiting further iteration ,as seen when CEM-R reaches an ER of 54.4% in round three without replay. CEM-R, with its larger data capacity, achieves significant performance improvements but shows the greatest decline in other tasks, likely due to the increased volume of CPT data diluting the replay effect. Combining CEM with other advanced replay-based methods warrants further investigation.
6 Discussion
Why can the Extractive Corpus be effective? LLMs using the Extractive Corpus show higher accuracy and W2R metrics. Unlike the Review Corpus, where the role in performance improvement is intuitive, the precise mechanism by which the Extractive Corpus augments the model’s capability to capture pivotal knowledge in lengthy texts remains worthy of discussion.
Extractive Corpus provides the model with both questions and supplemental material containing relevant knowledge. Questions and options typically present straightforwardly and can be condensed into several key points, while the supplemental material tends to intricately interweave numerous factual statements Jiang et al. (2024). Thus, we suggest that questions act as a structured induction of the supplemental material, where the Extractive Corpus explicitly links complex information to the resulting induction (i.e., the question and options). Questions and options can serve as cues and guides for the model to gather relevant information from the corpus, helping it consider how this knowledge is accessed through questions. This process enhances the model’s ability to extract useful information from the corpus while filtering out redundant and irrelevant details.
@@ 7 Conclusion
In this paper, we propose the Continue Evolving from Mistakes (CEM) method, a data-efficient and cost-effective method for collecting CPT data and continually evolving LLM on specific tasks by identifying and addressing internal knowledge deficiencies revealed by model’s mistakes. We believe that the difficulty of collecting CPT data and the limitations of instruction tuning’s superficial alignment have become a bottleneck for further performance improvement. To address this, we design a pipeline to collect corpus from mistakes and a novel parallel dataset construction paradigm using both CIT and CPT data to achieve continual and efficient internal knowledge supplementation during fine-tuning stage. We conduct experiments on several open-source LLMs to validate the efficacy of CEM. The results consistently showing significant improvements in answer accuracy and potential for continual evolution through multiple iterations.
@@ Appendix D Format of CIT Data
Table 3 and Table 4 show the IFT data formats of the Normative Corpus and the Extractive Corpus.
An Example of Normative Corpus
@@
Table 3: An example illustrating the format of the Normative Corpus.
An Example of Extractive Corpus
@@
Table 4: An example illustrating the format of the Extractive Corpus. The ellipsis indicates that some supplemental text has been omitted for clarity.
Appendix E Entailment Checking of Extractive Corpus
Table 5 shows the prompt used for entailment checking of the Extractive Corpus with Qwen2-72B-Instruct.
@@ Appendix E Entailment Checking of Extractive Corpus
Table 5: An example illustrating the entailment checking of the Extractive Corpus. The ellipsis indicates that some supplemental text has been omitted for clarity.
@@ Appendix F Exploring Fine-Tuning Strategies
Appendix G Data Division
Table 7 shows the the data division of datasets. For Xiezhi and CMMLU, the multiple-choice questions are divided into three parts. The Pre-Finetuning part is utilized to train the model to answer questions in a normative format before supplemental training. The Testing part is used to construct the test set, and the remaining part is used to construct the Normative Corpus and Extractive Corpus for supplemental training.
Dataset Total Pre-Finetuning Testing Supplemental Training Random Replay
Xiezhi 27,265 10,000 10,000 7,265 -
CMMLU 11,917 4,000 5,000 2,917 -
HotpotQA 7,000 2,000 5,000 - 500
GSM8K 7,000 2,000 5,000 - 500
Table 7: The data division and usage of Xiezhi, CMMLU, HotpotQA, and GSM8K. For HotpotQA and GSM8K, the data used for Random Replay is included in pre-finetuning.
Appendix H Experimental Result
Table 8 presents the comprehensive experimental results for various LLMs employing the CEM method on the Xiezhi and CMMLU tasks, with the supplemental training dataset constructed using the Plain Strategy.
Table 9 shows the complete experimental results of CuteGPT-13B on the Xiezhi task after being trained with supplemental training sets constructed using the Plain, Extractive, and Review strategies, as well as various data sources.
Table 10 presents the comprehensive experimental results of CuteGPT-13B on the Xiezhi task following two rounds of CEM, with the second round employing the same training data construction strategy as the first.
Table 11 presents the comprehensive results of catastrophic forgetting experiments for CuteGPT-13B on the Xiezhi, GSM8K, and HotpotQA tasks after one, two, and three rounds of CEM, respectively.
Task Metrics CuteGPT-13B Qwen-7B-Chat ChatGLM-6B
INIT Wiki Bing INIT Wiki Bing INIT Wiki Bing
Xiezhi Acc 42.93 45.91 46.56 41.62 48.71 47.87 40.44 51.95 53.61
W2R - 22.20 21.91 - 22.17 20.94 - 26.01 25.86
R2W - 18.57 18.93 - 15.08 14.70 - 14.50 12.69
CMMLU Acc 26.44 32.30 35.32 24.72 39.12 40.80 23.00 35.62 35.26
W2R - 23.84 25.14 - 25.24 26.94 - 25.96 26.32
R2W - 17.98 16.26 - 11.04 11.02 - 13.34 14.06
Table 8: Experimental results of various LLMs using the CEM method on the Xiezhi and CMMLU tasks
Data Metrics INIT CEM-N CEM-E CEM-E’ CEM-R0 CEM-R0.5 CEM-R1 CEM-R’
Wiki Acc 42.93 46.56 52.03 50.69 53.40 55.18 54.67 49.44
W2R - 22.20 24.12 23.42 24.82 24.93 24.61 22.90
R2W - 18.57 15.02 15.65 14.35 12.68 12.87 16.38
Bing Acc 42.93 45.91 48.47 45.11 53.90 53.86 54.53 46.58
W2R - 21.91 23.15 21.37 24.99 24.25 25.18 22.28
R2W - 18.93 17.61 19.18 14.02 13.32 13.58 18.62
Mix Acc 42.93 46.61 53.89 50.78 53.37 53.42 54.43 53.07
W2R - 21.99 25.12 23.67 25.00 24.04 25.06 24.81
R2W - 18.31 14.16 15.81 14.56 13.55 13.46 14.67
Table 9: The table shows the experimental results of CuteGPT-13B on the Xiezhi task using different supplemental training set construction methods.
Metrics Wiki Bing
CEM-N CEM-E CEM-R CEM-N CEM-E CEM-R
Acc INIT 42.93 42.93 42.93 42.93 42.93 42.93
Round1 46.56 52.02 55.18 45.91 48.46 53.86
Round2 51.69 55.20 59.93 46.37 48.74 58.33
W2R vs. INIT 24.46 26.03 28.29 22.07 23.32 27.01
vs. Round1 20.00 17.56 16.43 18.49 18.33 16.61
R2W vs. INIT 15.70 13.76 11.29 18.63 17.51 11.61
vs. Round1 14.87 14.39 11.68 18.03 18.06 12.14
Table 10: The table shows the experimental results of CuteGPT-13B using the CEM method on the Xiezhi task, where "v.s. INIT" indicates W2R and R2W calculated based on the initial evaluation results, and "v.s. Round1" indicates W2R and R2W calculated based on the evaluation results after one round of CEM.
Round Strategy Xiezhi GSM8K HotpotQA ER AFR
INIT 40.84 13.20 80.60 1.00 1.00
Round 1 CEM-P 42.02 6.64 51.60 1.03 0.730
CEM-E 42.76 8.36 55.52 1.05 0.681
CEM-R 51.49 7.54 55.62 1.26 0.673
Round 2 CEM-P 48.68 9.94 57.84 1.19 0.723
CEM-E 49.47 9.76 60.24 1.21 0.746
CEM-R 54.23 6.88 56.68 1.33 0.703
Round 3 CEM-P 50.58 10.02 60.96 1.24 0.757
CEM-E 52.69 9.46 62.98 1.29 0.772
CEM-R 58.91 7.48 59.86 1.44 0.678
Table 11: The table shows the experimental results of CuteGPT-13B using the CEM method on the Xiezhi, GSM8K, and HotpotQA tasks.

### v6

@@ 5.3 Mechanistic Analysis of Extractive and Review Instruction Components
Figure 5: The figure presents the R2W and W2R metrics of the models after CEM supplemental training.
@@ 5.4 Multiple Iterations Analysis
To evaluate the iterability of the CEM method and its potential for catastrophic forgetting during multiple training rounds, we conduct three rounds of CEM training on CuteGPT-13B-ift for the Xiezhi task, using consistent training strategies and non-repetitive CIT data of equivalent size in each round. Specifically, the CEM method is applied to the Xiezhi task, while HotpotQA and GSM8K are used to assess reasoning and mathematical skills. The ER metric evaluates performance improvements on Xiezhi, and the AFR metric monitors performance degradation in other tasks. Initially, 2,000 instances from each dataset are sampled for pre-finetuning to establish a baseline. To address forgetting, we implement a random replay strategy, replacing General Instructions with 500 randomly replayed instances from HotpotQA and GSM8K in each training round.
Figure 6: The figure shows the experimental results of CuteGPT-13B-ift on the Xiezhi task after multiple rounds of CEM. The left part represents results without replay, while the right part shows results with a random replay strategy. ER indicates the enhancement rate on the Xiezhi task, and AFR represents the average forgetting rate on the HotpotQA and GSM8K tasks.
@@
Figure 6 shows that even after three rounds of CEM, performance on Xiezhi improves significantly and maintains an upward trend. The combination of CEM and random replay effectively mitigates performance decline, with AFR exceeding 75.0% in most cases, highlighting its iterative potential. However, without a replay strategy, initial rounds show better improvement (possibly because the model does not need to adapt to the more complex multitasking required by random replay), but more rounds lead to greater forgetting, limiting further iteration ,as seen when CEM-R reaches an ER of 62.4% in round three without replay. CEM-R, with its larger data capacity, achieves significant performance improvements but shows the greatest decline in other tasks, likely due to the increased volume of CPT data diluting the replay effect. Combining CEM with other advanced replay-based methods warrants further investigation.
@@ 6 Conclusion
In this paper, we propose the Continue Evolving from Mistakes (CEM) method, a data-efficient and cost-effective method for collecting CPT data and continually evolving LLM on specific tasks by identifying and addressing internal knowledge deficiencies revealed by model’s mistakes. CEM employs a pipeline that leverages model mistakes to gather extensive CPT data, combined with a novel parallel training set construction paradigm integrating both CIT and CPT data for continuous training. Experiments on various open-source LLMs show that CEM significantly improves answer accuracy and demonstrates strong potential for sustained evolution through multiple iterations.
@@ Appendix D Format of CIT Data
Table 4 and Table 5 show the IFT data formats of the Normative Instruction and the Extractive Instruction.
An Example of Normative Instruction
@@
Table 4: An example illustrating the format of the Normative Instruction.
An Example of Extractive Instruction
@@
Table 5: An example illustrating the format of the Extractive Instruction. The ellipsis indicates that some supplemental text has been omitted for clarity.
Appendix E Entailment Checking of Extractive Instruction
Table 6 shows the prompt used for entailment checking of the Extractive Instruction with Qwen2-72B-Instruct.
@@ Appendix E Entailment Checking of Extractive Instruction
Table 6: An example illustrating the entailment checking of the Extractive Instruction. The ellipsis indicates that some supplemental text has been omitted for clarity.